I need this with work shown

5. Solve the equation for x: (3.x - 1)2 + 6(3.1 - 1) - 7= 0.
Answer #1

Kindly go through the solution and let me know in case of any doubt or further clarification is needed in the comment box.
Thanks for the question :)
3x-126(3x-) -1-0 (3x-1 3x) - Éx-1)-7 = o Ex-1 (3x-1)( 3x -1+1) = 0 3xt6) (3x-1-1) = 0 (3x+6 (3x-2) = 3x 6 - X -2 2, 3